OpenStudy (anonymous):

6 workers can construct a kitchen in 17 hours. How long should it take 9 workers to construct the kitchen?

OpenStudy (anonymous):

11.22 hours

OpenStudy (anonymous):

6 workers - 17 hours to do work. => 1 worker does 1/6 th the work in 17 hours. => 1 worker does the work in 17*6 hours. Therefore if there are nine people each person does 1/9 th the work. Time taken = 1/9 th the time taken by one worker to complete work = 17*6 /9 = 11.22

OpenStudy (anonymous):

\[\frac{17}{6}=\frac{x}{9}\] \[x=\frac{17\times 9}{6}\]

OpenStudy (anonymous):

six is two thirds of nine. so two third of seventeen is 11.22